# Learning on LLM Output Signatures for Gray-Box Behavior Analysis

### v2

@@ B.6 Ablation Study
The Role of the TDS ($\mathbf{X}$). As a case study, we examine both the DCD task on the BookMIA dataset and the HD task across the three datasets: HotpotQA, IMDB, and Movies. Figures 6 and 7 report a close-up comparison between LOS-Net and our two proposed baselines, which explicitly neglect the TDS, namely, ATP+R-Transf. and ATP+R-MLP. These plots consistently show how the best-performing model is LOS-Net. In many cases, LOS-Net outperforms the alternatives by a significant margin, indicating that the information encoded in the TDS ($\mathbf{X}$) is crucial for both tasks. Regarding the two ATP-based baselines, we report that ATP+R-Transf. obtains better performance than ATP+R-MLP in 8 out of 12 cases, but these improvements do not seem to follow a clear pattern across LLMs and datasets. The only exception is BookMIA, on which the former architecture outperformed the latter across all the four attacked LLMs.
Figure 8: Ablation study analyzing the effect of the hyperparameter $K$ introduced in Equation 1.
The hyperparameter $K$. To evaluate the impact of the hyperparameter $K$ introduced in Equation 1, we conduct a comprehensive case study focusing on the task of HD.
We experiment with various values of $K$, specifically $K\in\{10,50,100,500,1000\}$, and trained the same selected model whose results are reported in Table 1 for $K=1000$. The corresponding Test AUCs are presented in Figure 8.
@@
B.7 Results For Hallucination Detection for Qwen-2.5-7b
Table 4 reports results on our three considered HD datasets over LLM Qwen-2.5-7b-Instruct (Q-2.5-7b) [53]. We can see LOS-Netoutperforms all non-learnable output-based baselines by large margin, as well as our learnable baselines ATP+R-Transf. and ATP+R-MLP.
B.8 Results On The WikiMIA Dataset
@@
[…]
As shown in Table 5, LOS-Net consistently surpasses all baseline methods across all eight combinations of LLMs and datasets. Notably, for L-30b, our model achieves an AUC score that is more than 8 points higher than the best-performing baseline, MinK%++ for both datasets, demonstrating a substantial improvement. Similarly, for P-6.9b, our model maintains a steady advantage of approximately 5 AUC for both datasets, further underscoring its robustness. Overall, the second-best method is MinK%++, followed by MinK%, consistent with the findings of [58].
B.9 LOS-Net Run-Time
In Table 6, we report the wall-clock training times (for the best selected model based on the held-out validation set) and single-example detection times for LOS-Net  for all experiments presented in this paper.
@@ Appendix C Additional Tasks Background
Large-scale pre-training of LLMs typically involves crawling vast amounts of online data, a common practice to meet their substantial data requirements. However, this approach risks exposing models to evaluation datasets, potentially compromising our ability to assess their generalization performance accurately [6], or, taking a different perspective, can pose legal and ethical issues when models are accidentally exposed to copyrighted or sensitive data during training. This phenomenon is typically referred to as Data Contamination. Recently, Li et al. [26] demonstrated that LLMs from the widely used LLaMA [46] and Mistral [21] model families exhibit significant data contamination, particularly concerning frequently used evaluation datasets.
Hallucination Detection.
@@
LLMs’ tendency to generate untrustworthy outputs, commonly known as "hallucinations," remains a significant challenge to their widespread adoption in real-world applications [45]. To address this issue, various hallucination mitigation techniques have been proposed, including retrieval-augmented generation [24, 19, 12], customized fine-tuning [33, 7, 37], and, inference-time manipulation [25, 38, 59], to name a few. However, applying these methods to all user-LLM interactions can be computationally expensive. As a more targeted approach, hallucination detection has been explored to enable selective intervention only when necessary.
General Considerations on Annotations.
@@
We consider access to a set of annotations $y$’s, which we naturally associate with the corresponding LOS elements. These encode ground-truth labels pertaining to problems of interest, e.g., whether the input text $\vec{s}$ is in the pretraining corpus of $f$, or whether $f$ hallucinated when generating $\vec{g}$ from prompt $\vec{s}$. Collecting these annotations is generally possible, and various strategies could be adopted. For example, for DCD, labels can be gathered with collaborative efforts testing for text memorization, as studied e.g. in [11]. We also note that annotations are immediately (and trivially) available for open-source LLMs with disclosed pretraining corpora such as Pythia [4]. As we demonstrated in Section 5, models trained on annotations available for one LLM can, in some cases, be transferred and applied to another LLM.
@@
For HD, ground-truth labels can be collected by providing the target LLM with inputs prompting for completion or question answering on known facts and/or reasoning tasks. Hallucinations or error annotations are derived by comparing the consistency of the model’s response with known, factually true, or logically correct answers. For further details, refer to Section B.5.1.
Appendix D Broader Impact
By improving the detection of data contamination and hallucinations in Large Language Models, our contribution could foster the responsible development and use of Generative AI by enhancing transparency and trustworthiness. Applications to detecting machine-generated content could also support reducing the spread of AI-fabricated misinformation. We also note, however, that our work sheds light on predictive information contained in the output next-token probability distributions. This could lead malicious actors to develop more sophisticated defense mechanisms to MIAs or lead companies to potentially limit access to LLM outputs, thus hindering advancements in open research.
Appendix E LOS-Net Visualization
In Figure 9 we provide a visualization of our architecture, LOS-Net .

### v3

@@ B.7 Ablation Study
The Role of the TDS ($\mathbf{X}$). As a case study, we examine both the DCD task on the BookMIA dataset and the HD task across the three datasets: HotpotQA, IMDB, and Movies. Figures˜6 and 7 report a close-up comparison between LOS-Net and our two proposed baselines, which explicitly neglect the TDS, namely, ATP+R-Transf. and ATP+R-MLP. These plots consistently show how the best-performing model is LOS-Net. In many cases, LOS-Net outperforms the alternatives by a significant margin, indicating that the information encoded in the TDS ($\mathbf{X}$) is crucial for both tasks. Regarding the two ATP-based baselines, we report that ATP+R-Transf. obtains better performance than ATP+R-MLP in 8 out of 12 cases, but these improvements do not seem to follow a clear pattern across LLMs and datasets. The only exception is BookMIA, on which the former architecture outperformed the latter across all the four attacked LLMs.
Figure 8: Ablation study analyzing the effect of the hyperparameter $K$ introduced in Equation˜1.
Figure 9: Probability mass ($y$-axis) as a function of $K$ ($x$-axis) for each LLM/dataset combination considered in the HD study.
The hyperparameter $K$. To evaluate the impact of the hyperparameter $K$ introduced in Equation˜1, we conduct a comprehensive case study focusing on the task of HD.
We experiment with various values of $K$, specifically $K\in\{10,50,100,500,1000\}$, and trained the same selected model whose results are reported in Table˜1 for $K=1000$. The corresponding Test AUCs are presented in Figure˜8.
@@
To complement Figure˜8, we present Figure˜9, which shows the average probability mass captured for each value of $K$, computed as $\nicefrac{{\sum_{n=1}^{N}\left(\sum_{v=1}^{V}(\mathbf{X}^{\prime}_{n,v})\right)}}{{N}}$. Across all LLM/dataset combinations and for every $K\in\{10,50,100,500,1000\}$, the captured probability mass exceeds 91%. This helps explain why even small values such as $K=10$ perform well, as observed in Figure˜8.
@@
B.8 Results For Hallucination Detection for Qwen-2.5-7b
Table˜5 reports results on our three considered HD datasets over LLM Qwen-2.5-7b-Instruct (Q-2.5-7b) [53]. We can see LOS-Net outperforms all non-learnable output-based baselines by large margin, as well as our learnable baselines ATP+R-Transf. and ATP+R-MLP.
B.9 Results On The WikiMIA Dataset
@@
The WikiMIA-32 and -64 datasets contain excerpts from Wikipedia articles, consisting of, resp., 32 and 64 words. The distinction between contaminated and uncontaminated data is determined by timestamps. As in [43, 58], we attack Mamba-1.4b [15] (M-1.4b), LlaMa-13b/30b [47] (L-13b/30b), Pythia-6.9b [5] (P-6.9b).
Since WikiMIA does not provide an official training split and our method requires labeled data, we perform 5-fold cross-validation with training, validation, and testing splits [^15] We use $\{\frac{3}{5},\frac{1}{5},\frac{1}{5}\}$ as the ratios for training, validation, and testing, respectively. and rerun all baselines under the same protocol for a fair comparison. Results are reported as the mean and standard deviation across folds. For these datasets only, setting the hyperparameter $K=1000$ (recall Equation˜1) led to suboptimal performance in preliminary experiments, thus, we set $K=\text{``Full-Vocabulary''}$.
As shown in Table˜6, LOS-Net consistently surpasses all baseline methods across all eight combinations of LLMs and datasets. Notably, for L-30b, our model achieves an AUC score that is more than 8 points higher than the best-performing baseline, MinK%++ for both datasets, demonstrating a substantial improvement. Similarly, for P-6.9b, our model maintains a steady advantage of approximately 5 AUC for both datasets, further underscoring its robustness. Overall, the second-best method is MinK%++, followed by MinK%, consistent with the findings of [58].
B.10 Empirical Run-Times
In Table˜7, we report the wall-clock training times (for the best selected model based on the held-out validation set) and single-example detection times for LOS-Net   for all experiments presented in this paper.
@@ Appendix C Additional Tasks Background
Data Contamination Detection..  Large-scale pre-training of LLMs typically involves crawling vast amounts of online data, a common practice to meet their substantial data requirements. However, this approach risks exposing models to evaluation datasets, potentially compromising our ability to assess their generalization performance accurately [7], or, taking a different perspective, can pose legal and ethical issues when models are accidentally exposed to copyrighted or sensitive data during training. This phenomenon is typically referred to as Data Contamination. Recently, Li et al. [28] demonstrated that LLMs from the widely used LLaMA [47] and Mistral [23] model families exhibit significant data contamination, particularly concerning frequently used evaluation datasets.
@@
Hallucination Detection..  LLMs’ tendency to generate untrustworthy outputs, commonly known as "hallucinations," remains a significant challenge to their widespread adoption in real-world applications [46]. To address this issue, various hallucination mitigation techniques have been proposed, including retrieval-augmented generation [26, 21, 14], customized fine-tuning [34, 8, 39], and, inference-time manipulation [27, 40, 59], to name a few. However, applying these methods to all user-LLM interactions can be computationally expensive. As a more targeted approach, hallucination detection has been explored to enable selective intervention only when necessary.
@@
General Considerations on Annotations..  We consider access to a set of annotations $y$’s, which we naturally associate with the corresponding LOS elements. These encode ground-truth labels pertaining to problems of interest, e.g., whether the input text $\vec{s}$ is in the pretraining corpus of $f$, or whether $f$ hallucinated when generating $\vec{g}$ from prompt $\vec{s}$. Collecting these annotations is generally possible, and various strategies could be adopted. For example, for DCD, labels can be gathered with collaborative efforts testing for text memorization, as studied e.g. in [12]. We also note that annotations are immediately (and trivially) available for open-source LLMs with disclosed pretraining corpora such as Pythia [5]. As we demonstrated in Section˜5, models trained on annotations available for one LLM can, in some cases, be transferred and applied to another LLM.
@@
For HD, ground-truth labels can be collected by providing the target LLM with inputs prompting for completion or question answering on known facts and/or reasoning tasks. Hallucinations or error annotations are derived by comparing the consistency of the model’s response with known, factually true, or logically correct answers. For further details, refer to Section˜B.5.1.
Appendix D LOS-Net Visualization
In Figure˜10 we provide a visualization of our architecture, LOS-Net .
@@ Appendix D LOS-Net Visualization
Appendix E LLM Processing Pipeline
In Figure˜11 we provide a visualization of the LLM processing pipeline.
Appendix F Importance of TDS Illustration
We demonstrate the importance of the TDS tensor through the following example, see Figure˜12
Figure 12: Illustrative example of the importance of the TDS.